The midpoints of square ABCD are connected to form square EFGH. If EF=10units,
what is the area of squareABCD? Could u solve this please?

Answer by venugopalramana(3286)

The midpoints of square ABCD are connected to form square EFGH. If EF=10units,
what is the area of squareABCD? Could u solve this please?
JOIN AC...A DIAGONAL OF THE SQUARE.
IN TRIANGLE ABC,WE HAVE E IS THE MIDPOINT OF AB AND F IS THE MIDPOINT OF BC.
SO EF IS EQUAL TO HALF OF AC
EF=10
AC=2*10=20....SIMILARLY IN THE SQUARE THE OTHER DIAGONAL BD=20
AREA OF A SQUARE =PRODUCT OF DIAGONALS/2=20*20/2=200 SQUARE UNITS.